FEL Upgrade July 2002 Monthly Report


A major milestone was achieved this month with the delivery of the second cryomodule into the FEL vault. Transport of the module was uneventful. By the end of the month the module was positioned correctly and then cooled down to 2K. Kudos to the SRF, cryo and installation groups plus our own I&C team for their support of this important milestone.

We were involved in DOE’s annual review of Jefferson Lab’s Science and Technology programs July 19-21.

Our first dipole magnet for the FEL Upgrade, one of the new injector magnets (GV), was delivered by the end of the month. This magnet will be the first new dipole to go through the magnetic measurement for qualification and then onto installation in the injector area.

The I&C team has successfully demonstrated a new video capture system that will allow us to automate electron beam brightness (i.e., emittance) measurements in the FEL Upgrade driver accelerator.

We are pleased to congratulate Wael Ibrahim who completed his requirements for a PhD degree in Electrical Engineering at Old Dominion University on the subject of "Ultrafast electron dynamics in metals". Wael did work using the FEL under the guidance of Michelle Shinn and his thesis advisor, Prof. Hani El Sayed-Ali, a charter member of our Laser Processing Consortium.


We are saddened by the sudden death of one of our skilled and dedicated colleagues over the July 4th weekend. Among other duties, Karel Capek was our builder, assembler and installer of diagnostic equipment into the electron beam lines. Without him, there is a huge void in the team of steady handed folks that build the machine and keep it running. We will miss him.

Fred Dylla, Stan Majewski, Andrei Afanasev, Cynthia Keppel of Hampton U., and Alan Todd of AES attended the Workshop on the Role of Nuclear Physics Research Community in Combating Terrorism in Washington, July 11-12. The meeting was sponsored by the DoE Division of Nuclear Physics. We presented information on accelerator applications, nuclear particle detection, THz imaging, and the production of anti-microbial surfaces.

George Neil attended a briefing for Dr. Sega, Chief Technologist in the Office of the Secretary of Defense on accelerating funding for the Navy High Power FEL scaleup plan. Also attending were a number of high level military and civilians including Adm. Mathis, and Adm. Cohen. Approval was given to proceed to the next level. On Monday July 19th, Fred Dylla and George Neil participated in a briefing to the Vice Chief of Naval Operations, Admiral James Fallon given by Admiral Mathis of NAVSEA and Admiral Cohen of ONR concerning the status and plans for developing FELs for Naval applications. A follow-up briefing to the CNO is being scheduled for next month. (This was subsequently waived by the CNO and replaced by scheduled visits to the Office of the Secretary of the Navy).

On July 19-21, the science and technology of the FEL project were reviewed as part of DOE’s annual review of the lab’s S&T programs. The review team gave complimentary remarks on all of the lab’s programs and as expected provided advice on how to continue to optimize productivity when resources are constrained. With respect to the FEL program the reviewers were "impressed with the accomplishments of the FEL team given the relatively short life of the program" and also complimented the FEL user program start-up, which although very limited in resources, was "productive" with noteworthy accomplishments in nanotube production, ablations studies and protein dynamics. The first significant demonstration of an "energy recovered linac" with IR Demo was deemed very important for future machine development for both the nuclear science and light source communities. The synergy of the FEL program with the nuclear physics program at JLab is characterized as "excellent" with the cautionary note that potential conflicts should be avoided or minimized by continued vigilance of management.

The Upgrade project monthly report for the month of June 2002 was completed and distributed to the DOE, ONR and AFRL program managers.

Funding ($567k) for the two JTO projects that were funded this year on: (1) injector drive laser development and (2) high power cryogenic optical cavity mirrors was received by our DOE Site Office this month. We thank our contract monitors at JTO, ONR and DOE for negotiating the bureaucratic path that got the funding to us.

We spent some time analyzing the calculations that are used to support the ODH (oxygen deficiency hazard) classification of the FEL Facility prior to the cool down of the cryomodules on Wednesday July 25. A Temporary Safety Operating Procedure (TSOP) was put in place to allow the installation and cool down operations to proceed until the final configuration of passive and active controls are in place.

Discussions were held with our Engineering Department to estimate our engineering support needs for the remainder of this fiscal year and the next fiscal year for final installation and commissioning activities, in addition to new design tasks associated with the JTO projects and any potential start-up of 100 kW FEL hardware.

On July 24-25, we had visitors from AES, Inc for preparation of our presentations to the JTO- Navy review of injector systems for high power FELs which was held at AES’s Medford, NY facility on July 31-Aug.1.

We presented a summary at this review of experience with the IR Demo photogun which operated at 5mA, our improvements incorporated in the 10 mA FEL Upgrade gun, and a summary of our collaboration with AES on the development of a 100 mA gun for the proposed 100 kW FEL Upgrade.

Project Cost Performance:

The Phase 1 FEL Upgrade project budget for the period June 1, 2000 to September 30, 2001 is $9,029k. The first phase of the project was fully obligated as of the end of September. The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion.

The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002. Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k. Work for Phase 2 proceeded to plan during this reporting period. Actual charges of $384k were accrued for July for total accrued costs of $4,021k to date for Phase 2. Work scheduled for July was $129k for a total of $3,518k for Phase 2 to date; work performed for July was $420k for a total of $3,515k for Phase 2 to date, which results in 79% scheduled vs. 79% performed for the Phase 2 project. At present the cost variance for Phase 2 is -$507k and the schedule variance is -$4k. We expect to meet the planned cost at completion of Phase 2 effort. The program is presently on track to begin commissioning efforts in September 2002.

The UV FEL project was approved for $2,836k for a performance period from September 2001 to September 30, 2002. Actual charges of $138k were accrued through July for a total accrued costs of $1,944k since project start. Work scheduled for July was $434k for a total of $2,135k for the UV project to date, work performed for July was $190k for a total of $1,892k for the UV project to date, which results in 75% scheduled vs. 67% performed for the UV project. At present the cost variance for the UV project is -$51k and the schedule variance is -$243k. We had to slow down the efforts on the UV project in July because of the late delivery of FY02 funds from the Air Force. These funds were not available until August 2. The program is on track to meet the statement of work requirements on schedule and within the available budget.

WBS 3 (Beam Physics):

A "People's Injector Subcommittee" (Benson, Hernandez, Douglas) of the commissioning politburo met to discuss injector recovery. Most parameters are well characterized using Demo procedures; a PARMELA study is planned to develop means to lock in the few remaining indeterminants (buncher phase & gradient, SRF cavity 4 phase).


WBS 4 (Injector):

The ball cathode was implanted successfully with the W&M implantation system. The system is now fitted with two roughing pumps, making the plasma significantly more. The gun stand and chamber were aligned and the SF6 tank was connected so positioning of the high voltage tank could be done.

The tapered shielded bellows and shielded / BPM bellows drawings were signed off and submitted to the shop for manufacturing.

Continued to work with W&M on the implantation of a test tube. Two runs were completed successfully at half the nominal operating voltage to minimize the possibilities of arcing. No arcs occurred during the runs and the implanted layer is strongly attached to the surface of the tube despite the lower operating voltage. The SF6 tank modification drawings were signed off and submitted to the shop. The rear 6-way cross and cathode shield operating bellows were submitted to the shop for final welding.

The light box can is complete. The machine shop is now working on completing the mirror holders that insert into the can. The installation crew has installed two of the four stands for the drive laser transport line.

The door for the HVPS Tank was hung this week. The mating flanges, o-ring and grove are being cleaned and greased.


Preparations continued for the 8-cavity string assembly for the first CEBAF Upgrade cryomodule.

Cavity parts and components are being assembled for the second upgrade module which will be assembled next spring for the 3rd FEL cryomodule.

WBS 6 (RF):

Quarter Klystrons - LCW was plumbed onto the klystron cart for unit 3 this month. Nearly all interlocks have been wired and tested.

Quarter HVPS - The last of the interlocks are being wired and tested this month. Unit 3 will be ready for high voltage next week. The SOP is being circulated comments.

Zone 4 - Requisitions were written and signed today to purchase the flexible waveguides necessary to complete the installation to the second cryomodule. Several additional sections of waveguide will also need to be procured.

WBS 8 (Instrumentation):

The BLM front card schematic is done. The drawing has been checked and approved. The pcb layout for the board is well under way and attention is being focused on procurement of parts and device package selection (details). The Fast Raster board check plots have been completed by EECAD and are currently being Q.A.'d for final sign off.

A design review was held for the new Machine Protection System (MPS) and the Drive Laser Pulse Controller (DLPC) this month. There are a couple of 'features' that need to be worked out; Steve Benson has the action item to work out with G. Neil and M. Shinn. The marching orders are now complete for the entire Machine Protection System.

All the magnet trim cables (0F through 5F regions) have been pulled to their respective destinations in the vault. Locations and channel assignments for the BLM's have been identified and signal and HV cables have been re-routed and pulled into place.



The 5 mm hole was successfully placed in the 3 linac beam viewer foils. There are 4 additional foils ready for mounting. The first 4 DV trim magnets were mounted to the girders and are having their input leads configured. The major installation activity for the week was the reworking of the video system. The video crosspoint is being upgraded to 192 video inputs to 16 outputs (these are blocks of 64X16). The field wiring to each of the beam viewer video is included in this effort. The VME video frame grabber software was upgraded, as soon as the video is back up (next week) the viewer calibration files will begin to be created.

We received the F0151 BLM front PCB layout (component placement) and routing has begun. This effort should take a week then we will go out for the four required boards (total 48 channels). We have begun an EPICS on Linux effort with 2 for the students. One of the applications is to use a PC/104 board as an IOC and put it in the tunnel to control the IR optical cavity motors and TC guages, the other is to do video processing of the beam viewer signals to have an automated emittance measurement.

The existing Drive Laser Pulse Controller (DLPC) has been put back together and is ready for FEL commissioning. The upgraded pulse controller will be brought online in parallel with the old system. The first component of the new DLPC has been implemented in iocfel10. It is an F0080 Timer Card which will be used for synthesizing the FEL's pulsed beam repetition rate and handle the beam synchronization requirements detailed in the design review. The DLPC design review presentation can be viewed by clicking link below and a screen capture of the expert EPICS screen for the DLPC's Rep-Rate Timer card as well as a screen-shot from the oscilloscope showing the results can be viewed by clicking the link below. The design for the C0305 (micropulse frequency generator) carrier board is almost finished. *DLPC design review->http://laser/techdoc/FEL_Project_Review/revans/DLPC_2002_A.htm *DLPC screen-shots> http://www.jlab.org/ccc/mail_archives/ACCELERATOR/fel_ctrls/CURRENT/msg00003.html. All parts are on order to finish-up the F0151 (VME based BLM cards) project. The F0151 layout is nearing completion and is expected to go out for production next week. Additionally, the software to utilize the new hardware is also well in progress.

Control system debug continued on the beamviewer system for the 3F region. Troubleshot and repaired beamviewer chassis' and control cabling. Video cables for the beamviewers have been pulled into position. Video rack (FL02B09) has been disassembled and reworked, hardware in the rack has been tested to accommodate the upgrade.

The new video capture system is up and running. This new system is based on a PC running Linux and acting as an IOC talking directly to EPICS. The PC and capture card are running under Linux now and the EPICS will come on soon. Pavel Evtushenko will be visiting in September and will be working on a fully incorporating this to an automated emittance measurement system. Additionally the PC104 system for control of the optical cavity is also making progress. Vellums for the Fast Raster boards have been received from EECAD and signed off. The raster boards have been sent to the board fabrication house with a 5-day turn around. Six Canbus interface boards were populated. These will be used for the new communications to the beam viewer control chassis. The Quad magnet wiring harnesses have been completed for the machine.

A great deal of progress has been made on DC power this month. There was a design review for the 100 Amp sextapole and octapole power supplies. The design will be based on a commercial switcher and an external current transducer with 5 ppm stability. An interface chassis will be built that contains a switch (for bipolar) and a closed loop regulator. The (deeper) racks for these supplies arrived, were installed, and had the three phase 208 plumbed in. The drawings are being finalized for the higher current cable runs. The 100Amp cable is on order and the 535MCM cable will be reused. The wiring harness's for the DB/DJ correctors have been fabricated and will be installed on the girders as they arrive in the vault. Termination of the magnet trim cables in junction boxes 4 and 5 has been completed, rung out and labeled. The conversion of the trim racks from CAMAC based controls to a serial interface is in progress.



Both the Switched Electrode Electronics (SEE) and the four channel Beam Position Monitor (BPM) cables were reinstalled this week. There are an additional 8 BPMs so more cable will be ordered. The RF modules for the SEE's have been turned over to the I&C group in the main machine for testing and calibration.

Effort continued on the HV interlock chassis. The design is complete and awaiting EECAD for

schematic and then PCB layout. The new design will combine functions from three different chassis from the IRDEMO

Progress also continues on learning how to efficiently use the Motorola 807 DSP. The application that is being worked is processing the signals from a 128-segment line scan camera and outputting the FEL spectrum (using a grating). We had started with few ADC programming examples available from Motorola, to understand the working of on-board ADCs. There are 2 options. While talking to the board, firstly generating code using Motorola provided drivers and libraries (called the SDK tools). Secondly generating code without using Motorola drivers or libraries. (1)The previous modules (presently running them as separate programs) developed for Clkin_gen and Beamsync_Pulse(o/p) have now been combined and run using SDK. (2)Rewritting the programs separately without using the SDK tools now. This would ensure that the execution would be faster and code more readable by the user if changes are required.

Lastly the Video acquisition system is up and running under Linux. The software is being worked on to create the tools for an automated emittance measurement system. We will have the PC running Linux acting as an IOC with no VXWorks license fees. The automatic calibration (magnification) and background subtraction routine are being coded now.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. the all GW dipole had their Purcell Gap shims successfully glued down. Variation (with respect to a surface plate) over the entire active area of each pole is of the order of .001 inch. This variation is the combination of two epoxy bonds and two sheets of material plus the flatness of the pole. This is excellent work. In addition to booties and hairnets and clean room, and tack rags, three gluing folks at MEI shaved their heads to make sure no fallen hair got into the glue joints.

• The units were painted, reassembled and shipped

• Cores for the Second batch of 6 Magnets for UV are almost complete.

• The epoxy from the formulator that is with out flexibilizer was used on coils 14 through 18 and they have no surface welts. This cosmetic problem is solved and the solution will be used at Wang NMR for the other impregnated coils.

Injector Dipoles (DU/DV)

• At Wang NMR the first three GVs were put together. They discovered a few missing machined and hardware parts. Some problems were solved by Master Machine shipping the missing items in the above shipment or by Wang purchasing hardware from local suppliers.

• Near the end of the month, Wang NMR discovered they are still missing two of the 24 field clamp parts in the package from Master Machine. Finally, the missing GV field clamp parts were found at Master Machine, sent to Wang and became part of the assembly

• The first GV was shipped by airfreight.

• The two GU cores were received by Wang NMR underwent trial assembly.

• Wang NMR glued the shim materials to the some GU cores.

Arc 180 Degree Dipoles (GY)

• At Wang NMR, the third and fourth GY Coils were wound.

• Wang has been doing quality checks in anticipation of potting the first three coils. The first coil has not been able to be made short free because of some built-in twist to the conductors formed during the winding process. They believe the coil is not reparable in a reliable way and should be scrapped. The remainder of the coils test out good. They have enough conductor to wind the first two of the four layers of a fifth coil. We ordered more conductor this week to complete the coil. We are resolving the contractual problems resulting from having to wind a fifth coil.

• GY Coil Potting fixture remains in fabrication.

• Three GG coils were potted with the flexiblized epoxy and look good. Wang NMR will pot the remainder with the non-flexiblized epoxy. Arrangements are being made to exchange epoxy with P. D. George Co., the formulator.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) completed one of the GQ Cores. Gap dimensions were inspected by me on July 11 and passed. Again, extraordinary tolerances (.0003 in.) were held by another manufacturer.

• The shim pieces for the energy change were water jet cut to rough size through 3.5 inches of steel producing thin slices with no distortion. These pre heat-treated slices will be ground to final size using a vacuum chuck.

• At Wang NMR, all eight coils for the GQ Dipoles are wound and are being prepared for potting and the GX coil winding is almost complete. We discussed potting fixture construction and methods. They will use vacuum pressure impregnation with the non flexiblized epoxy.

• Process Equipment Co. (PECo) has checked the gap of the three remaining GQ cores and is doing the final machining on them to get the tooling balls in the right locations.

• The shim pieces for the energy change were are ground to size.

• Process Equipment Co. (PECo) assembled the remaining three GQ cores and will ship early next month. The GX cores are on track for delivery two weeks later.

• Wang NMR received 10 gal. of the alternate (non-welting) epoxy formulation from P D. George Co.


3 inch quad (QX)

• Four more girders were installed on their stands. More girders were assembled, adding to a stockpile that will be installed when personnel are available. The installers have been assigned to the CEBAF during its month-long shut down.

Trim Quad (QT)

• Milhous Control of Virginia finally got two production units to us.

• We are now analyzing the use of the integral GC vertical corrector at several levels of quadrupole operation in anticipation of the change request required for this innovation

• We have run the first unit and have found that the side coils are not as well cooled as we would like. The first unit passed temperature test with added cooling of the side coils. We generated an engineering change order with the details so that Milhous can add this feature to the existing units.

Sextupole (SF)

• We finished checking the sextupole drawings from DULY Research. We were disappointed with the quality of the drawings and requested some rework at their expense. We also made substantial improvements to the intra coil connection design (from discussions with MEI) using wrought fittings and the cover. All the corrections were sent to DULY to start the drawing rework.

• DULY research will finish the Sextupoles over the next two weeks.

Octupole (OT)

• DULY started design of the Octupole.

Beam Line and Vacuum

• Master Machine started machining the flanges of the Arc Vacuum Chambers and received the material. They are working with a vendor with a 5 axis water jet cutting machine to cut out the parts. We inspected a sample piece of stainless steel that had been water jet cut and determined that there was minimal contamination from the cut method and that in addition, the style of the weld joint would lead to minimal contamination of the vacuum surfaces. We concluded that water jet cutting for these chambers is acceptable.

• The JLab Shop started tack welding the X Chamber together and received material to start fabrication of the four chambers for the optical chicane region.

• Design of the regions around the wiggler is 70 % complete

• The drawings of the four chambers for the optical chicane region were signed off for Jlab Shop fabrication. The materials have already been purchased.

• The stands for sextupoles and trim quads are in procurement.

• Design of the Octupole stand started.

• The first meeting was held to start design of the recirculation dump region.


WBS 11 (Optics):

Both optical cavity assemblies arrived this month and were installed on the test & assembly stands. This gives us the opportunity to try various combinations of lead screw/nut combinations to see which has the lowest friction (Rigorous particulate generation checks will probably wait until we can mount them in the vacuum vessels). We received the final QC on the OCMMS beamline parts. Several small variances (e.g., two planes parallel to 0.003" instead of 0.002") from the drawings were found. We are meeting today to formally review the results and sign off on them. The OCMMS support stands were received. We have firm quotes on the optical breadboards and are awaiting release of funds to order them. The first prototype, compact 50 kW beam dump is brazed. We are getting quotes to have it coated. We received two bids on the optical cavity vacuum vessels. Both were higher than our preliminary estimates, so we are going back to the low bidder to see what drives the price. We anticipate making an award next week. We have a designer working on the insertable mirrors. Excellent progress is being made, and we will hold a preliminary design review in about a week. We are in the process of developing a LabVIEW program to speed measurements of outgassing rates of optical cavity components.

We have quotes for a new grating that extends the IR range of our spectrometer, and will purchase it. We have been confirming the performance of the ultrafast laser oscillator (amplitude jitter, power & pointing stability, etc) and so far it is exceeding specifications. We have one more measurement (timing jitter) which we'll measure today. The first test of our FROG (FEL pulseshape diagnostic) setup is going well. We received our new IR beam profiler (a Spiricon PyroCAM III) which will improve our capabilities.

This month’s tests with the optical cavity hardware turned up another problem with the rotary feedthrough - ~ 10 degrees of backlash before the shaft turned. Discussions with the manufacturer suggest they know the problem existed but it wasn't published in their literature. What concern us are the consequences of the feedthrough internals failing and rupturing the bellows, and thus venting the chamber. We are revisiting and considering other vendors to find if a better solution exists.

Problems with the motion control software/hardware continue to be worked. We found that an overnight, ~ 100 deg. C bake of a cleaned and disassembled LVDT resulted in a pressure of 3E-8Torr. This is close to the level of performance we need for the IR machine, so we will continue working with this vendor to reduce the outgassing. Tests of the OCMMS beamline hardware show that the mechanical tolerancing was more than adequate. We are in the process of finding the best compromise between ROC measurement accuracy and beam vignetting on the upstream assembly, caused by the requirement to use a 2" dia BPM, rather than the 3" variety.

Our tests of the prototype backplane-cooled mirror assembly in the mirror test stand showed that the design does a good job of dissipating thermal loading higher than anticipated. The total wavefront aberration at the highest loading we tried, 65 W/sqr. cm. was ~ 1 wave at 633 nm. The aberration is linear with the absorbed power, and at the specified power output and loading, is only ~ 0.2 waves. Since we wish to make the optical transport readily upgradeable to higher FEL outputs, this suggests we make the substrate thinner, and we will discuss this with various vendors to see if that is feasible.

Progress on the UV upgrade:

We've received additional optics for the UV optics damage test project. We are setting up the hardware, and are awaiting sign off of the safety documents. We are continuing to survey the literature for the most recent absorption measurements of UV optics. As the IR optical transport system (OTS) design progresses, we are looking ahead at incorporating these design elements into the UV OTS design.


Aerospace/Jefferson Lab Subcontract

Several aspects of the engineering design were investigated in further detail. The proposed system will use two electro-optic Pockel (EOP) cells in series for controlling the FEL laser beam power and pulse repetition rate. The EOP will operate at FEL powers of less than 5 Watts. Discussions with the vendor showed that particular design of EOP could be so configured if the polarization of the incident laser was of high quality. We are still discussing means to control the EOPs with signals generated from motion controller software.

The software subsystem was evaluated with several vendors for capabilities to control >10 axes of motion and be scalable with additional dependent and independent axes. Our proposed design has at least 3 axes that are synchronous are dependent. Discussions with the vendors showed that a special virtual axes can be developed that could be slaved to the three main axes and provide information on the vector velocity of the XYZ system.

The XYZ motion control hardware vendor conducted additional tests for increasing the patterning speed while maintaining an position error resolution to less than 200 nm. Several different driver amplifier schemes were attempted on various precision stages. The current data show that it will be possible to pattern microstructures at speeds approaching 20mm/sec and maintaining position error resolution of better than 200nm.

The general overall design of the laser microengineering station was drawn to scale using a solid modeling software. Before the design is finalized models will have to be evaluated in terms of maintaining flatness and isolating vibration.

The design team has agreed to submit a report on the current design. This is in addition to the presentation material to be made available for the systems requirement review, now scheduled for August 22, 2002.


Other Related Projects


IR Upgrade Commissioning (ONR)


This project was funded in FY02 with $1,000k, with a period of performance from July 1, 2002 to March 30, 2003. In this reporting period, $260k has been expended on this project for a total expenditure to date of $260k.


Work was initiated on a detailed commissioning plan for use when driver accelerator commissioning begins on or about October 1, 2002. The initial focus is on the start-up of the injector. Pre-commissioning activities continued on the RF systems. The Zone 4 (CM2) RF system is ready for SRF commissioning after installation of the final waveguide connections to cryomodule 2. Interlocks have been tested on the injector klystron stands. With regard to pre-commissioning activities for the beam transport sytem, the magnet test stand for the dipoles was readied for the impending receipt of the injector dipoles (DU, DV) next month. In preparation for injector commissioning, the existing drive laser pulse controller (DLPC) was re-commissioned while the upgraded DLPC is completed in parallel.


Cryomodule 3


This project was funded in FY02 with $400k. Additional funding of $350k is anticipated in FY03. In this reporting period, $25.8k was expended on this project, for a total expenditure to date of $25.8k.


Cavities and cavity end groups are being assembled for the 3rd cryomodule. Design and fabrication of the tooling required for the cryomodule assembly is proceeding. This latter activity is being cost-shared with the production of the first CEBAF-Upgrade cryomodule, which will precede the FEL 3rd cryomodule.

One Micron Wiggler Project (ONR)


This project was funded in FY02 with $600k. Completion funding of $600k is anticipated in FY03. In this reporting period, $34.1k has been expended on this project for a total expenditure to date of $34.1k.


A study was carried out to compare the optical klystron and simple uniform wiggler options. A spreadsheet was developed to calculate optical klystron gain and power vs. laser and electron beam parameters. The spreadsheet model for a uniform wiggler was checked against pulse propagation simulation results. The gain was in good agreement and the power was found to be low. Initial results indicate that the gain for a given power output is smaller for the optical klystron and the Rayleigh range is larger resulting in a factor of two higher intensity on the mirrors for the optical klystron.

JTO Cryomirror Project


Funding for FY02 was received July 16, 2002. A budget of $286.4k is available for the effort. We initiated work on the program and are making efforts to recover lost time.


During this reporting period expenditures of $19.2k were incurred. Commitments total $25.8k. Total expenditures to date are $19.2k. The program is on track financially and is expected to deliver all elements of the SOW within projected cost.


Initial work centered on modeling the deformation of the cryomirror under the expected heat load. A finite element model of the mirror was established and initial indications are that maximum temperature rise for a sapphire mirror initially at 40 K is 0.8 K at 10 W absorbed and 8.7 K for 100 W absorbed. This level of absorbed power brackets the 50 W that would correspond to 1 MW FEL output with a cavity outcoupling of 20%. During the next reporting period we intend to study thermal shock effects on the mirror substrate and mounting by cooling a mockup system repeatedly in liquid nitrogen. From the modeling effort, we will calculate the maximum deformation due to the imprint of the Gaussian beam and the resulting change in the radius of curvature of the cavity mirror.


JTO Drive Laser Project


Funding for FY02 was received July 16, 2002. A budget of $264.1k is available for the Phase 1 effort. We initiated work on the program and are making efforts to recover lost time.


During this reporting period expenditures of $7.8k were incurred. Total expenditures to date are $ 7.8k. The program is on track financially and is expected to deliver all elements of the SOW within projected cost.


Procurement specifications for the master oscillator front end system are in preparation. A preferred technical approach has been determined and we are going to request quotes from several vendors. We expect to begin procurement in August.